40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 14 23:14:24 2018, aaron, Update, OMC, OMC scanning/aligning script 
    Reply  Tue Nov 20 20:33:11 2018, aaron, Update, OMC, OMC scanning/aligning script pathwaysketch.pdfIMG_0094.JPGIMG_0097.JPG
       Reply  Mon Nov 26 15:43:16 2018, aaron, Update, OMC, OMC scanning/aligning script c1omc.png
          Reply  Thu Dec 6 11:16:28 2018, aaron, Update, OMC, OMC channels 
             Reply  Mon Dec 10 12:29:05 2018, aaron, Update, OMC, OMC channels 52DE723A-02A4-4C62-879B-7B0070AE8A00.jpeg545E5512-D003-408B-9F00-55F985966A16.jpegDFF34976-CC49-4E4F-BFD1-A197E2072A32.jpeg
                Reply  Mon Dec 10 19:47:06 2018, aaron, Update, OMC, OMC channels 
                   Reply  Tue Dec 11 13:42:44 2018, Koji, Update, OMC, OMC channels 
                   Reply  Tue Dec 11 13:48:04 2018, aaron, Update, OMC, OMC channels 
                      Reply  Thu Dec 13 22:24:21 2018, aaron, Update, OMC, OMC channels 
          Reply  Mon Dec 10 12:11:28 2018, aaron, Update, OMC, Aligning the OMC B9DCF55F-1355-410C-8A29-EE45D43A56A4.jpeg
             Reply  Tue Dec 11 14:24:18 2018, aaron, Update, OMC, Aligning the OMC 
                Reply  Tue Dec 11 22:50:07 2018, aaron, Update, OMC, Aligning the OMC 6x
                   Reply  Thu Dec 13 22:36:42 2018, aaron, Update, OMC, Aligning the OMC 
                      Reply  Fri Dec 14 13:02:24 2018, aaron, Update, OMC, Aligning the OMC 
                      Reply  Fri Dec 14 13:06:12 2018, aaron, Update, OMC, Aligning the OMC 
                         Reply  Mon Dec 17 20:45:40 2018, aaron, Update, OMC, Aligning the OMC IMG_5922.JPGIMG_5921.JPG
                            Reply  Tue Dec 18 18:13:32 2018, aaron, Update, OMC, OMC L HV Piezo driver tests (again) IMG_4576_copy.jpg
                               Reply  Wed Dec 19 00:12:46 2018, gautam, Update, OMC, 40m OMC DCC node 
Message ID: 14343     Entry time: Tue Dec 11 14:24:18 2018     In reply to: 14337     Reply to this: 14346
Author: aaron 
Type: Update 
Category: OMC 
Subject: Aligning the OMC 

I set up a function generator to drive OMC-L, and have the two DCPD mons and the OMC REFL PD sent to an oscilloscope. I need to select a cds channel over which to read the REFL signal.

The two DCPD mon channels have very different behaviors on the PD mons at the sat box (see attachment). PD1 has an obvious periodicity, PD2 has less noise overall and looks more white. I don't yet understand this, and whether it is caused by real light, something at the PDs, or something at the sat box.

I've again gone through the operations that will happen with the OMC chamber vented. Here's how it'll go, with some of the open questions that I'm discussing with Gautam or whoever is around the 40m:

  1. Function generator is driving OMC-L. Right now there is one 150V Kepco supply in use, located on the ground just to the right of the OMC rack. I only have plans to power it on while scanning OMC-L, and until the OMC is fully in use the standard practice will be to use this HV with two people in the lab and shut it off after the immediate activities.
    1. To do: Is a second drive necessary for the TT drivers? I don't think it is during this vent, because we will want to align into the OMC with the TTs in a 'neutral' state. I recall that the way the TT drivers are set up, 0V from the dac to the driver is the 'centered' position for all TTs. Unless we want to compensate for some known shift of the chambers during pumpdown, I think this is the TT position we should use while aligning the OMMT into the OMC.
    2. To do: make sure I'm driving the right pins with the function generator. Update: Seems I was driving the right channels, here's the pinout.
  2. We will use the reflection of aux from the SRM to align into the OMC.
    1. Gautam pointed out that I hadn't accounted for the recombination BS for the aux beam being 90-10. This means there's actually something like 300uW of aux onto the OMC, rather than ~3mW. This should still be enough to see on a card, so it is fine.
    2. However, the aux beam is aligned to be colinear with the AS beam when the SRM is misaligned. So the question is whether the wedge on the SRM makes the SRM-reflected aux beam not colinear with the AS beam

 

---------

Talked with Gautam for a good while about the above plan. In trying to figure out why the DCPD sat box appears to have a different TF for the two PDs (seems to be some loose cabling problem at the mons, because wiggling the cables changed this), we determined that the AA chassis also wasn't behaving as expected--driving the expected channels (28-31) with a sine wave yields some signal at the 100Hz driving frequency, but all save ch31 were noisy. We also still saw the 100Hz when the chassis was unplugged. I will continue pursuing this, but in the meantime I'm making an IDE40 to DB37 connector so I can drive the ADC channels directly with the DAC channels I've defined (need to match pinouts for D080303 to D080302). I also will make a new SCSI to DB37 adapter that is more robust than mentioned here. I also need to replace the cable carrying HV to the OMC-L driver, so that it doesn't have a wire-to-wire solder joint.

We moved a razor blade on the AP table so it is no longer blocking the aux beam. We checked the alignment of aux into the AS port. AUX and AS are not colinear anywhere on the AP table, and despite confirming that the main AS beam is still being reflected off of the OMC input mirror, the returning AUX beam does not reach the AP table (and probably is not reaching the OMC). AUX needs to be realigned such that it is colinear with the AS beam. It would be good if in this configuration, the SRM is held close to its position when the interferometer is locked, but the TTs should provide us some (~2.5mrad) actuation. Gautam will do this alignment and I will calculate whether the TTs will be able to compensate for any misalignment of the SRM.

Here is the new plan and minimal things to do for the door opening tomorrow:

  1. Function generator is driving OMC-L.
    1. The PZT mon channel is sent to the oscilloscope.
    2. To do: confirm again that the triangle wave I send in results in the expected triangle wave going to the OMC, using this mon channel.
  2. The OMC REFL signal is being sent to the AP PD. See photo.
    1. Need to align into this PD, but this alignment can be done in air on the AP table.
  3. Monitor the DCPD signals using the TPs from the sat box going to the oscilloscope.
    1. There may be further problems with the sat box, but for the initial alignment into the OMC only the REFL signal is necessary.
    2. Not minimally necessary, but the sat box needs a new case. It has a front, back, and bottom, but no main case, so the board is exposed.
  4. I will move the OMMT-to-OMC steering mirrors while watching the scope for flashes in the REFL signal.

That is the first, minimal sequence of steps, which I plan to complete tomorrow. After aligned into the OMC, the alignment into the DCPDs shouldn't need modification. Barring work needed to align from OMC to DCPDs, I think most other work with the OMC can be done in-air.

ELOG V3.1.3-